# thanks: Equal contributionthanks: Equal contribution

### v1

@@ Abstract
Ecological interactions can dramatically alter evolutionary outcomes in complex communities. Yet, the classic theoretical results of population genetics (e.g., Kimura’s fixation formula) largely ignore ecological effects. Here, we address this shortcoming by using dynamical mean-field theory to integrate ecology into classical population genetics models. We show that ecological interactions between parents and mutants result in frequency-dependent selection and can be characterized by a single emergent parameter that measures the strength of ecological feedbacks. We derive an explicit analytic expression for the fixation probability that generalizes Kimura’s famous result and analyze it in various ecological and evolutionary limits. We find that ecological interactions suppress fixation probabilities for moderately beneficial mutants when compared to Kimura’s predictions, with the strength of suppression increasing with larger effective population sizes, greater selective differences between parent and mutant, and for ecosystems with a large number of “open niches” (i.e., ecosystems well below the packing bound). Frequency-dependent selection also gives rise to prolonged parent-mutant coexistence in complex communities, a phenomenon absent in classical population genetics. Our study establishes a framework for integrating ecological interactions into population genetics models and helps illuminate how ecology can affect evolutionary outcomes.
@@ Introduction
Virtually all models of population genetics make a simplifying assumption: that the evolutionary dynamics of mutants are determined only by their interaction with their parents. In doing so, they neglect a fundamental biological fact: almost all evolving populations are embedded in complex ecological communities with which they interact [14, 15, 16]. These ecological interactions can fundamentally alter evolutionary trajectories [17, 18, 19]. Any comprehensive theory of evolution must therefore incorporate ecological interactions. Without such an eco-evolutionary theory, we cannot understand evolution in natural contexts [20, 21, 22].
@@
Numerous works have sought to address these limitations. These include models with frequency-dependent selection that capture more realistic evolutionary scenarios [23, 24, 25, 26]. However, in these models the relationship to ecology remains implicit and unspecified [27, 28]. Other evolutionary theories seek to integrate ecological context [29, 30] but many of these efforts remain qualitative in nature and give rise to very different predictions [31, 32, 33, 34, 35, 36]. More quantitative approaches often make particular assumptions about the nature of ecological interactions that make it hard to assess their general applicability. This includes assuming low community diversity [37, 38], specific fitness functions [39] and specialized ecological models [40, 18]. Further, many of these approaches ignore stochasticity due to small mutant population sizes [40, 18], which crucially sets the “drift barrier” that mutants must cross to fix. In short, we still currently lack a general theory that can predict and explain the dynamics of mutants in highly-diverse communities across a wide variety of ecological models.
@@
At first, incorporating all the complex details of community ecological interactions might seem intractable due to the large number of unknown parameters. However, recent approaches from statistical physics and random matrix theory have revealed that for a sufficiently complex ecological community, it is possible to make strong predictions about ecological properties of an ecosystem such as its diversity or stability [41, 42, 43, 44, 45, 46, 47, 48]. A key insight from these works has been the realization that at the community level, the effects of ecological interactions can often be summarized in terms of a few emergent parameters that relate to susceptibilities that measure the strength of ecological feedback [49]. Inspired by these efforts, we may use tools from statistical physics to generalize Kimura’s theory to incorporate ecological interactions.
@@
Here we derive an effective theory of population genetics in complex ecological communities. We use dynamical mean-field theory (DMFT) to coarse-grain the effect of a highly-diverse community on a mutant’s frequency dynamics. Our central result is that a community’s combined effect can be captured through a single emergent parameter that quantifies the strength of community-mediated feedbacks on mutant dynamics. This parameter can be computed from a community’s statistical properties such as its niche-packing. We demonstrate that community-mediated feedbacks generically lead to frequency-dependent selection across a variety of ecological models. This can dramatically alter the fixation probabilities of mutants compared to classical population genetic predictions. Further, in diverse communities, mutants can often coexist with their parents for extremely long periods, which drastically alters their fixation and extinction times. Our results highlight that the fate of a mutant can be quantitatively predicted even in complex ecological settings.
@@
Figure 1: Population genetics and parent-mutant dynamics in the absence and presence of complex communities. (a) Classic population genetics typically models dynamics only for the parent and the mutant, independent of community context. (b) In complex communities, ecological interactions among species collectively influence the parent-mutant dynamics. As shown using dynamical mean-field theory, the ecological community mediates an effective frequency-depdendent selection between the parent and the mutant. (c) Parent-mutant dynamics can be expressed in terms of mutant frequency $f$. (d) In classical population genetics, the dynamics of $f$ are driven by constant selection and stochastic drift. (e) In complex communities, additional frequency-dependent selection arises, with an emergent parameter $\eta$ characterizing the strength of ecological feedbacks. (f-i) Examples of dynamics of species abundances in classical population genetics (left) versus in complex communities (other species’ dynamics in grey). After a long time, the mutant eventually reaches either extinction (f–g) or fixation (h–i).
@@ Results
Ecological feedbacks manifest as frequency-dependent selection. Population genetics aims to describe how mutations spread through a species population. To do so, it follows the dynamics of a mutant’s frequency $f(t)$ in the population given by
@@
Classical population genetics makes a simplifying but crucial assumption: the selection coefficient $s$ remains constant throughout the evolutionary process (Fig. 1d) [1, 2, 4], though see exceptions [11]. This assumption effectively treats parents and mutants as if they are isolated from interactions with an ecological community. For this reason, it neglects all the ways in which ecological communities can potentially alter evolutionary trajectories through environmental feedbacks (Fig. 1b) [14, 15, 16]. This assumption becomes especially hard to justify in complex communities which are highly-diverse and ubiquitous in nature: from rainforests to microbiomes.
@@
with the $\{\xi_{i}(t)\}$ independent normal random variables with $\langle\xi_{i}(t)\rangle=0$ and $\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. In this expression, $r_{i}$ is the carrying capacity of species $i$; $A_{ij}$ captures ecological interactions between species and encodes how species $j$ affects the abundance of species $i$; and $\sqrt{2DN_{i}}\xi_{i}(t)$ represents demographic noise. The parameter $D$ sets the overall scale of demographic noise and serves as an analog of a temperature or diffusion coefficient. Following recent work, we model parameters $r_{i}$ and $A_{ij}$ as random variables [43, 46, 51] drawn from Gaussian distributions (Methods).
To study population genetics in the GLV, we initialize a community with $S$ randomly chosen species at steady-state (Methods). We then pick one of these species at random to be the parent $p$. We denote the abundance of the parent species in the community by $N_{p}(t)$. At time $t=0$, we introduce a mutant $m$ to this community at a small abundance $N_{m}(0)\ll N_{p}(0)$. This is equivalent to assuming that the mutant starts at low frequency $f_{0}\ll 1$. To model the relatedness of parent and mutant, we assume that the mutant’s interaction parameters, $A_{mj}$, with other species in the community are correlated with those of the parent, $A_{pj}$. We denote the strength of this correlation as $\rho$.
Consistent with population genetics, we also assume that mutants and parents interact much more strongly with each other than with any other species in the community. To model this, we set parent-mutant interactions $A_{mp}=A_{pm}=\rho\approx 1$ while interactions with all other species in the community are assumed to be much weaker ($\approx 1/S$, Methods). As we will show below, even though mutants interact weakly with any individual species in the community, these weak effects add up due to the large number of species ($\approx S$) present in the community. Thus as a collective, community interactions can no longer be neglected and can strongly influence mutant dynamics. Throughout this work, we focus on the strong-selection-weak-mutation regime [52]: no subsequent mutations occur before a mutant fixes or goes extinct. For this reason, we ignore the possibility of multiple mutants.
@@
$\frac{df}{dt}=\left(s-\eta(f-f_{0})\right)f(1-f)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t),$ (3)
where $s$ measures the difference in invasion fitness between the mutant and parent into the community and plays a role analogous to selective advantage in classical population genetics, $N_{\text{eff}}=(N_{p}(0)+N_{m}(0))/2D$ measures the effective population size of the combined parent-mutant population, and $\xi(t)$ is a random normal variable with $\langle\xi(t)\rangle=0$ and $\langle\xi(t)\xi(t^{\prime})\rangle=\delta(t-t^{\prime})$.
This equation results from two distinct eco-evolutionary processes: deterministic selection – the term proportional to f(1f)f(1-f) – and stochastic drift – the term proportional to $\xi(t)$. Remarkably, just a single parameter $\eta$ suffices to capture the collective effect of the entire community on parent-mutant dynamics. Eq. (3) shows that selection is now frequency-dependent since the selection coefficient $s-\eta(f-f_{0})$ changes with mutant frequency. When $\eta=0$, e.g., in the absence of a community, we recover classical population genetics with a constant selection coefficient. However, when $\eta\neq 0$, ecological feedbacks fundamentally alter evolutionary dynamics.
We show that community-mediated feedbacks generically manifest as frequency-dependent selection across a wide variety of ecological models. In the SI, we derive expressions for the effective parameters—$s$, $N_{\text{eff}}$ and $\eta$—for MacArthur Consumer-Resource Models (CRMs) [54, 55] and generalized CRMs with nonlinear per-capita growth rates, which includes models with cross-feeding [56] (SI Sec. B and C). Remarkably, despite the diverse mathematical structures of these models, they all reduce to the same form of frequency-dependent selection. Thus, Eq. (3) may be interpreted as a universal effective model of population genetics in complex communities [5], analogous to Kimura’s model in classical population genetics (Fig. 1d).
@@
Starting with Eq. (3), we solved the corresponding backward equation to obtain an analytic formula for the fixation probability $p_{\mathrm{fix}}$ in complex communities (SI Sec. D):
$p_{\mathrm{fix}}=\frac{\mathrm{Erfi}\left(\alpha(\tilde{s}+f_{0})\right)-\mathrm{Erfi}\left(\alpha\tilde{s}\right)}{\mathrm{Erfi}\left(\alpha(\tilde{s}+f_{0})\right)-\mathrm{Erfi}\left(\alpha(\tilde{s}-1+f_{0})\right)}\,$, (5)
@@
$\tilde{s}=\frac{s}{\eta}\,$, (8)
@@
is the mutant-parent fitness difference normalized by the strength of the ecological feedback. The quantity $\tilde{s}$ is proportional to the “dressed invasion fitness” which was recently introduced in Ref. [17] for predicting the outcomes of ecological invasions. Eq. (5) serves as the generalization of Kimura’s formula in the context of complex ecological communities.
@@
To gain intuition for this formula, it is useful to look at various limits of this expression. When ecological effects are weak, i.e., $\alpha\ll 1$, our formula reduces back to Kimura’s original formula in Eq. (4) (SI Sec. D). To numerically confirm this, we simulated the dynamics of complex communities governed by the GLV model into each of which we introduced a mutant of a randomly chosen parent strain. We made ecological effects weak simply by setting a large value of $D$, increasing the strength of stochastic drift. We repeated simulations for mutants with a given invasion fitness difference $s$ between them and their parent and measured the fraction of simulations in which mutants fixed (Methods). As shown in Fig. 2a, the simulated fixation probabilities match our generalized formula remarkably well across a range of $s$. Further, in this limit, our formula virtually overlaps Kimura’s formula. In the next section, we will look at the opposite case of strong ecological effects, i.e., $\alpha\gg 1$, and show that we get a qualitatively different picture of the fate of a mutant compared to Kimura’s formula.
Figure 2: Fixation probabilities suppressed by strong ecological effects. We compute fixation probabilities $p_{\mathrm{fix}}$ at various invasion fitness differences $s$ by simulating mutations within generalized Lotka-Volterra models. Error bars denote standard errors from multiple instances of demographic noise and mutants. (a) When ecological effects are weak relative to stochastic drift ($\alpha\ll 1$), our theory matches Kimura’s formula for fixation probabilities and predicts them accurately in simulations. (b) When ecological effects are strong ($\alpha\gg 1$), Kimura’s formula applies only for deleterious or highly beneficial mutants. For moderately beneficial mutants with $s\lesssim\eta/2$, fixation probabilities are strongly suppressed relative to Kimura’s prediction. Our theory with ecology can capture such suppression and matches simulations. We give analytic expressions for fixation probability in (c) Kimura’s theory and (d) our theory with ecology. Ecological suppression of mutants depends on several factors: (e) For neutral mutations ($s=0$), we quantify the suppression by the percentage decrease in fixation probability from Kimura’s prediction ($1-p_{\mathrm{fix}}/f_{0}$). Suppression increases with larger effective population size $N_{\text{eff}}$ and lower parent-mutant correlation $\rho$. (f) Suppression is also stronger when the community is less packed (has more “open niches”), i.e., the number of surviving species is low relative to the bound set by competitive exclusion.
@@
Ecological interactions strongly suppress mutant fixation. In the presence of strong ecological effects, i.e., when $\alpha\gg 1$, our numerical simulations for the fixation probability show a stark deviation from Kimura’s predictions (Fig. 2b), with an almost switch-like behavior with increasing $s$. This is in contrast with Kimura’s formula which shows a gradual, linear increase in fixation probability over a range of $s$. The deviation from Kimura is most pronounced for moderately beneficial mutants where ecological effects strongly suppress the fixation probability (Fig. 2b, blue region). The origin of this suppression is easiest to understand for neutral mutants ($s=0$) where our formula reduces to (SI Sec. D)
$p_{\mathrm{fix}}(s=0)=\begin{cases}f_{0}&\alpha\ll 1\\ e^{-\alpha^{2}}&\alpha\gg 1\end{cases}\,.$ (9)
@@
Figure 3: Parent-mutant coexistence in complex communities. We compute mean absorption and extinction times at various invasion fitness differences $s$ by simulating mutations within generalized Lotka-Volterra models. Error bars denote standard errors from multiple instances of demographic noise. (a, b) The time scales of parent-mutant dynamics at various invasion fitness differences $s$ reveal the possibility of coexistence. Under strong ecological effects, the mean times to absorption (either extinction or fixation) (a) and to extinction (b) are exponentially longer for moderately beneficial mutants with $0\lesssim s\lesssim\eta$, peaking at $s\simeq\eta/2$ (dashed lines) as predicted by our theory. The maximum mean times also grow exponentially with the effective population size $N_{\mathrm{eff}}$. (c-h) Representative trajectories and phase portraits (in terms of the ratio between selection and drift) show distinct behaviors across different $s$. For low $s$ (f) or high $s$ (h), there is a single crossover (dashed lines) from drift to selection dominated regimes. Selection drives (green arrows) the mutant to extinction (c) or fixation (e). However, there are additional crossovers (g) when $s$ is moderate with $s\lesssim\eta$. Selection instead drives the mutant into a coexistence region centered at frequency $f^{*}$ (dark red), resulting in exponentially long times for parent-mutant coexistence (d).
@@
Using Eq. (5), we can also compute the full range of invasion fitness differences $s$ for which mutants are suppressed when ecological effects are strong (SI Sec. D). We find that suppression occurs when
$0\leq s\lesssim s_{c}:=\eta\left(\frac{1}{2}-f_{0}\right)\simeq\frac{\eta}{2}\,$, (10)
@@
where $s_{c}=\eta/2$ is the critical invasion fitness difference at which fixation probability switches from roughly 0 to almost 1. As we show in SI Sec. D, in this case the dominant contribution to ecological suppression in the range $0\lesssim s\lesssim s_{c}$ comes from the last term in the denominator of Eq. (5).
@@
The ecological suppression of mutants is an emergent community-mediated phenomenon whose strength depends on $\alpha=\sqrt{N_{\text{eff}}\eta}$. To understand how different ecological properties of the underlying community control the strength of suppression, we measured $\alpha$ across a variety of parameter sets in the GLV model and plotted the corresponding deviation of the simulated fixation probability from Kimura’s predictions for a neutral mutation ($s=0$) (see Fig. 2e). Suppression increased with the effective population size $N_{\mathrm{eff}}$ (Fig. 2e). This can be understood by noting that increasing $N_{\mathrm{eff}}$ weakens stochastic drift, and hence increases the importance of ecological feedbacks. Suppression also increased with decreasing parent-mutant correlation $\rho$ (Fig. 2e). The reason for this is that the ecological feedback parameter $\eta$, and hence $\alpha$, is proportional to $1-\rho$ (SI Sec. AC). Intuitively, as parent and mutant become less similar (i.e., $\rho$ decreases), mutants interact less strongly with their parents, allowing community feedbacks to play a stronger role in the dynamics.
@@
Parent-mutant coexistence alters fixation and extinction times. In addition to fixation probability, the second important quantity in population genetics is the time it takes for a mutant to reach fixation or extinction (which we refer to as absorption time). This quantity characterizes the timescale of adaptation in evolutionary dynamics. It is also informative to study the establishment time or equivalently the time to extinction. These two time scales are intimately related since extinction probability becomes negligible as soon as the mutant becomes established in the population, long before fixation. The mean extinction time can also be theoretically estimated (SI Sec. G). These times are relatively short in classic population genetics. Namely, they reach a maximum of order $N_{\mathrm{eff}}f_{0}|\log f_{0}|$ [58] for neutral mutations ($s=0$), and decay with the magnitude of the fitness difference $|s|$.

### v2

@@ Abstract
Ecological interactions can dramatically alter evolutionary outcomes in complex communities. Yet, the framework of population genetics largely neglects interactions from a species-rich community. Here, we bridge this gap by using dynamical mean-field theory to integrate community ecology into classical population genetics models. We show that ecological interactions result in emergent frequency-dependent selection between parents and mutants, characterized by a single parameter measuring the strength of ecological feedbacks. This result generalizes classical population genetics models to highly diverse communities and enables predictions of mutation outcomes in these eco-evolutionary settings. We derive an analytic expression for fixation probability that extends Kimura’s formula and reveals that ecological interactions strongly suppress the fixation of moderately beneficial mutations. This suppression arises because frequency-dependent selection leads to prolonged coexistence between parent and mutant lineages, which acts as a barrier to fixation. The strength of these effects increases with effective population size and the number of open niches in the ecosystem. Our study establishes a framework for integrating ecological interactions into population genetics, showing that evolutionary outcomes can be predicted using simple models even in the presence of complex community feedbacks.
@@ Introduction
While successful, these models make a simplifying assumption: that the evolutionary dynamics of mutants are mainly determined by their interaction with their parents. In doing so, they incorporate ecological context implicitly rather than explicitly. This makes it hard to understand how ecological interactions will alter evolutionary trajectories [14, 15, 16], which is important since almost all evolving populations are embedded in complex ecological communities with which they interact [17, 18, 19]. Any comprehensive theory of evolution must therefore explain how ecological interactions with a complex community will affect mutation outcomes. Without such an eco-evolutionary perspective, our understanding of evolution in natural contexts will remain incomplete [20, 21, 22].
@@
Numerous works have sought to bridge this gap. These include population genetics models with frequency-dependent selection that capture more realistic evolutionary scenarios [23, 24, 25, 26]. However, since these models assume communities with one to two species, it is unclear whether these apply to more complex ecological settings [27, 28]. Other evolutionary theories seek to integrate ecological context [29, 30] but many of these efforts remain qualitative in nature and give rise to very different predictions [31, 32, 33, 34, 35, 36]. More quantitative approaches often make particular assumptions about the nature of ecological interactions that make it hard to assess their general applicability. This includes assuming low community diversity [37], specific fitness functions [38], and specific choices of ecological models [39, 40, 15]. In short, we still currently lack general understanding of how population genetics may predict and explain the dynamics of mutants in highly-diverse communities across a wide variety of ecological models.
@@
At first, incorporating all the complex details of community ecological interactions might seem intractable due to the large number of unknown parameters. However, recent approaches from statistical physics and random matrix theory have revealed that for a sufficiently complex ecological community, it is possible to make strong predictions about ecological properties of an ecosystem such as its diversity or stability [41, 42, 43, 44, 45, 46, 47, 48]. A key insight from these works has been the realization that at the community level, the effects of ecological interactions can often be summarized in terms of a few emergent parameters that relate to susceptibilities that measure the strength of ecological feedback [49]. Inspired by these efforts, we may use tools from statistical physics to reveal links between ecological interactions and population genetics models.
@@
Here we derive an effective theory of population genetics in complex ecological communities. We use dynamical mean-field theory (DMFT) to coarse-grain the effect of a highly-diverse community on a mutant’s frequency dynamics. Our central result is that community-mediated feedbacks generically lead to emergent frequency-dependent selection across a variety of ecological models. A community’s combined effect can be captured by only a single new parameter that quantifies the strength of community-mediated feedbacks on mutant dynamics. This parameter can be computed from a community’s statistical properties such as its niche-packing. As a result, population genetics models with frequency-dependent selection are effective in describing mutant dynamics also in complex communities. We further find that community effects dramatically alter the fixation probabilities of mutants compared to Kimura’s predictions. Further, in diverse communities, mutants can often coexist with their parents for extremely long periods, which drastically alters their fixation or extinction times. Our results highlight that the fate of a mutant can be quantitatively predicted by population genetics models even in complex ecological settings.
@@
Figure 1: Population genetics and parent-mutant dynamics in the absence and presence of complex communities. (a) Classic population genetics typically models dynamics only for the parent and the mutant, independent of community context. (b) In complex communities, ecological interactions among species collectively influence the parent-mutant dynamics. As shown using dynamical mean-field theory, the ecological community mediates an emergent frequency-depdendent selection between the parent and the mutant. (c) Parent-mutant dynamics can be expressed in terms of mutant frequency $f$. (d) In classical population genetics, the dynamics of $f$ are driven by constant selection and stochastic drift. (e) In complex communities, additional frequency-dependent selection arises, with an emergent parameter $\eta$ characterizing the strength of ecological feedbacks. (f-i) Examples of dynamics of species abundances in classical population genetics (left) versus in complex communities (other species’ dynamics in grey). After a long time, the mutant eventually reaches either extinction (f–g) or fixation (h–i).
@@ Results
Ecological feedbacks manifest as emergent frequency-dependent selection. Population genetics aims to describe how mutations spread through a species population. To do so, it follows the dynamics of a mutant’s frequency $f(t)$ in the population given by
@@
Classical population genetics makes a simplifying but crucial assumption: the selection coefficient $s$ remains constant throughout the evolutionary process (Fig. 1d) [1, 2, 4], though see exceptions [11]. This assumption implies that ecological contribution in $s$, if any, remains constant. Effectively, this treats parents and mutants as if they are isolated from interactions with the dynamics of an ecological community. For this reason, it often misses the fact that ecological communities can potentially alter evolutionary trajectories through dynamical environmental feedbacks (Fig. 1b) [17, 18, 19]. This assumption becomes especially hard to justify in complex communities which are highly-diverse and ubiquitous in nature: from rainforests to microbiomes.
@@
with the $\{\xi_{i}(t)\}$ independent normal random variables with $\langle\xi_{i}(t)\rangle=0$ and $\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. In this expression, $r_{i}$ is the carrying capacity of species $i$; $A_{ij}$ captures ecological interactions between species and encodes how species $j$ affects the abundance of species $i$; and $\sqrt{2DN_{i}}\xi_{i}(t)$ represents demographic noise. The parameter $D$ sets the overall scale of demographic noise and serves as an analog of a temperature or diffusion coefficient. Following recent work, we model parameters $r_{i}$ and $A_{ij}$ as random variables [43, 46, 51] drawn from Gaussian distributions (Methods), although our results remain valid for other distributions with finite mean and variance [52, 53].
To study population genetics in the GLV, we initialize a community with $S$ randomly chosen species at steady-state (Methods). We then pick one of these species at random to be the parent $p$. We denote the abundance of the parent species in the community by $N_{p}(t)$. At time $t=0$, we introduce a mutant $m$ to this community at a small abundance $N_{m}(0)\ll N_{p}(0)$. This is equivalent to assuming that the mutant starts at low frequency $f_{0}\ll 1$. To model the ecological similarity of parent and mutant, we assume that the mutant’s interaction parameters, $A_{mj}$, with other species in the community are correlated with those of the parent, $A_{pj}$. We denote the strength of this correlation as $\rho$, which is independent of any difference in carrying capacity between parent and mutant.
Consistent with population genetics, we assume that mutants and parents are highly correlated ($\rho\approx 1$) in their interactions with the community. With a high $\rho$, the total parent and mutant population behaves effectively as a single species from the community perspective, which crucially simplifies the ecological dynamics (SI Sec. F). Moreover, mutants and parents interact much more strongly with each other than with any other species in the community. To model this, we set parent-mutant interactions $A_{mp},A_{pm}$ to be close to but weaker than the intraspecific competition ($\approx 1$, SI Sec. A), while interactions with all other species in the community are assumed to be much weaker ($\approx 1/S$, SI Sec. A). However, as we will show below, even though mutants interact weakly with any individual species in the community, these weak effects add up due to the large number of species ($\approx S$) present in the community. Thus as a collective, community interactions can no longer be neglected and can strongly influence mutant dynamics. Throughout this work, we focus on the strong-selection-weak-mutation regime [54]: no subsequent mutations occur before a mutant fixes or goes extinct. For this reason, we focus on a single mutation and ignore the scenarios of multiple mutants (known as clonal interference) or mutation rates. We also assume that the community is at an ecological steady state when the mutant appears.
@@
$\frac{df}{dt}=\left(s_{\mathrm{inv}}-\eta f\right)f(1-f)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t),$ (3)
where $s_{\mathrm{inv}}$ measures the invasion fitness of the mutant in the community. Since $s_{\mathrm{inv}}$ contains both the intrinsic selection coefficient $s$ (e.g., differences in carrying capacities) and community effects (SI Sec. A), it plays a role analogous to the mutant’s selective coefficient in classical population genetics. $N_{\text{eff}}=\left<N_{p}(0)+N_{m}(0)\right>/2D$ measures the effective population size of the combined parent-mutant population, which is approximately constant on average, especially for mutants where $\rho\approx 1$ (SI Sec. F). $\xi(t)$ is a random normal variable with $\langle\xi(t)\rangle=0$ and $\langle\xi(t)\xi(t^{\prime})\rangle=\delta(t-t^{\prime})$.
This equation results from two distinct eco-evolutionary processes: deterministic selection – the term proportional to f(1f)f(1-f) – and stochastic drift – the term proportional to $\xi(t)$. Remarkably, just introducing a single new parameter $\eta$ suffices to capture the collective effect of the entire community on parent-mutant dynamics. Eq. (3) shows that selection is now frequency-dependent since the selection coefficient $s_{\mathrm{inv}}-\eta f$ changes with mutant frequency. When $\eta=0$, e.g., in the absence of a community, we recover classical population genetics with a constant selection coefficient. However, when $\eta\neq 0$, ecological feedbacks fundamentally alter evolutionary dynamics.
[…]
Figure 2: Fixation probabilities suppressed by strong ecological effects. We compute fixation probabilities $p_{\mathrm{fix}}$ at various invasion fitness $s_{\mathrm{inv}}$ by simulating mutations within generalized Lotka-Volterra models. Error bars denote standard errors from multiple instances of demographic noise and mutants. (a) When ecological effects are weak relative to stochastic drift ($\alpha\ll 1$), our theory matches Kimura’s formula with $s=s_{\mathrm{inv}}$ for fixation probabilities and predicts them accurately in simulations. (b) When ecological effects are strong ($\alpha\gg 1$), Kimura’s formula applies only for deleterious or highly beneficial mutants. For moderately beneficial mutants with $s_{\mathrm{inv}}\lesssim\eta/2$, fixation probabilities are strongly suppressed relative to Kimura’s prediction. Our theory with ecology can capture such suppression and matches simulations. We give analytic expressions for fixation probability in (c) Kimura’s theory and (d) our theory with ecology. Ecological suppression of mutants depends on several factors: (e) For neutral mutations ($s_{\mathrm{inv}}=0$), we quantify the suppression by the percentage decrease in fixation probability from Kimura’s prediction ($1-p_{\mathrm{fix}}/f_{0}$). Suppression increases with larger effective population size $N_{\text{eff}}$ and lower parent-mutant correlation $\rho$. (f) Suppression is also stronger when the community is less packed (has more “open niches”), i.e., the number of surviving species is low relative to the bound set by competitive exclusion.
@@
Starting with Eq. (3), we solved the corresponding backward equation to obtain an analytic formula for the fixation probability $p_{\mathrm{fix}}$ in complex communities (SI Sec. D, see also [62]):
$p_{\mathrm{fix}}=\frac{\mathrm{Erfi}\left(\alpha\tilde{s}_{\mathrm{inv}}\right)-\mathrm{Erfi}\left(\alpha(\tilde{s}_{\mathrm{inv}}-f_{0})\right)}{\mathrm{Erfi}\left(\alpha\tilde{s}_{\mathrm{inv}}\right)-\mathrm{Erfi}\left(\alpha(\tilde{s}_{\mathrm{inv}}-1)\right)}\,$, (5)
@@
$\tilde{s}_{\mathrm{inv}}=\frac{s_{\mathrm{inv}}}{\eta}\,$, (8)
@@
is the mutant invasion fitness normalized by the strength of the ecological feedback. The quantity $\tilde{s}_{\mathrm{inv}}$ is proportional to the “dressed invasion fitness” which was recently introduced in Ref. [14] for predicting the outcomes of ecological invasions. Eq. (5) serves as the generalization of Kimura’s formula in the context of complex ecological communities.
@@
To gain intuition for this formula, it is useful to look at various limits of this expression. When ecological effects are weak, i.e., $\alpha\ll 1$, our formula reduces back to Kimura’s original formula in Eq. (4) (SI Sec. D) with $s=s_{\mathrm{inv}}$. To numerically confirm this, we simulated the dynamics of complex communities governed by the GLV model into each of which we introduced a mutant of a randomly chosen parent strain. We made ecological effects weak simply by setting a large value of $D$, increasing the strength of stochastic drift. We repeated simulations for mutants with a given invasion fitness $s_{\mathrm{inv}}$ and measured the fraction of simulations in which mutants fixed (Methods). As shown in Fig. 2a, the simulated fixation probabilities match our generalized formula remarkably well across a range of $s_{\mathrm{inv}}$. Further, in this limit, our formula virtually overlaps Kimura’s formula. In the next section, we will look at the opposite case of strong ecological effects, i.e., $\alpha\gg 1$, and show that we get a qualitatively different picture of the fate of a mutant compared to Kimura’s formula.
@@
Ecological interactions strongly suppress mutant fixation. In the presence of strong ecological effects, i.e., when $\alpha\gg 1$, our numerical simulations for the fixation probability show a stark deviation from Kimura’s predictions (Fig. 2b), with an almost switch-like behavior with increasing $s_{\mathrm{inv}}$. This is in contrast with Kimura’s formula which shows a gradual, linear increase in fixation probability over a range of $s$. The deviation from Kimura is most pronounced for moderately beneficial mutants where ecological effects strongly suppress the fixation probability (Fig. 2b, blue region). The origin of this suppression is easiest to understand for neutral mutants ($s_{\mathrm{inv}}=0$) where our formula reduces to (SI Sec. D)
$p_{\mathrm{fix}}(s_{\mathrm{inv}}=0)=\begin{cases}f_{0}&\alpha\ll 1\\ e^{-\alpha^{2}}&\alpha\gg 1\end{cases}\,.$ (9)
@@
Figure 3: Parent-mutant coexistence in complex communities. We compute mean absorption and extinction times at various invasion fitness $s_{\mathrm{inv}}$ by simulating mutations within generalized Lotka-Volterra models. Error bars denote standard errors from multiple instances of demographic noise. (a, b) The time scales of parent-mutant dynamics at various invasion fitness $s_{\mathrm{inv}}$ reveal the possibility of coexistence. Under strong ecological effects, the mean times to absorption (either extinction or fixation) (a) and to extinction (b) are exponentially longer for moderately beneficial mutants with $0\lesssim s_{\mathrm{inv}}\lesssim\eta$, peaking at $s_{\mathrm{inv}}\simeq\eta/2$ (dashed lines) as predicted by our theory. The maximum mean times also grow exponentially with the effective population size $N_{\mathrm{eff}}$. (c-h) Representative trajectories and phase portraits (in terms of the ratio between selection and drift) show distinct behaviors across different $s_{\mathrm{inv}}$. For low $s_{\mathrm{inv}}$ (f) or high $s_{\mathrm{inv}}$ (h), there is a single crossover (dashed lines) from drift to selection dominated regimes. Selection drives (green arrows) the mutant to extinction (c) or fixation (e). However, there are additional crossovers (g) when $s_{\mathrm{inv}}$ is moderate with $s_{\mathrm{inv}}\lesssim\eta$. Selection instead drives the mutant into a coexistence region centered at frequency $f^{*}$ (dark red), resulting in exponentially long times for parent-mutant coexistence (d).
@@
Using Eq. (5), we can also compute the full range of invasion fitness $s_{\mathrm{inv}}$ for which mutants are suppressed when ecological effects are strong (SI Sec. D). We find that suppression occurs when
$0\lesssim s_{\mathrm{inv}}\lesssim s_{c}:=\frac{\eta}{2}\,$, (10)
@@
where $s_{c}=\eta/2$ is the critical invasion fitness at which fixation probability switches from roughly 0 to almost 1. As we show in SI Sec. D, in this case the dominant contribution to ecological suppression in the range $0\lesssim s_{\mathrm{inv}}\lesssim s_{c}$ comes from the last term in the denominator of Eq. (5).
@@
The ecological suppression of mutants is an emergent community-mediated phenomenon whose strength depends on $\alpha=\sqrt{N_{\text{eff}}\eta}$. To understand how different ecological properties of the underlying community control the strength of suppression, we measured $\alpha$ across a variety of parameter sets in the GLV model and plotted the corresponding deviation of the simulated fixation probability from Kimura’s predictions for a neutral mutation ($s_{\mathrm{inv}}=0$) (see Fig. 2e). Suppression increased with the effective population size $N_{\mathrm{eff}}$ (Fig. 2e). This can be understood by noting that increasing $N_{\mathrm{eff}}$ weakens stochastic drift, and hence increases the importance of ecological feedbacks. Suppression also increased with decreasing parent-mutant correlation $\rho$ (Fig. 2e). The reason for this is that the ecological feedback parameter $\eta$, and hence $\alpha$, is proportional to $1-\rho$ (SI Sec. AC). Intuitively, as parent and mutant become less similar (i.e., $\rho$ decreases), mutants interact less strongly with their parents, allowing community feedbacks to play a stronger role in the dynamics.
@@
Parent-mutant coexistence alters fixation and extinction times. In addition to fixation probability, the second important quantity in population genetics is the time it takes for a mutant to reach fixation or extinction (which we refer to as absorption time). This quantity characterizes the timescale of adaptation in evolutionary dynamics. It is also informative to study the establishment time or equivalently the time to extinction. These two time scales are intimately related since extinction probability becomes negligible as soon as the mutant becomes established in the population, long before fixation. The mean extinction time can also be theoretically estimated (SI Sec. H). These times are relatively short in classic population genetics. Namely, they reach a maximum of order $N_{\mathrm{eff}}f_{0}|\log f_{0}|$ [64] for neutral mutations ($s=0$), and decay with the magnitude of the fitness difference $|s|$.